# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v2

@@ 4.2.4 Time Cost Comparison
4.3 Summary of Evaluation
@@
5 Conclusions
This paper presents TraCS-D and TraCS-C, providing local differential privacy guarantees for trajectory collection in continuous location spaces. TraCS-D uses a novel direction-distance perturbation procedure, while TraCS-C perturbs the Cartesian coordinates of each location. These two methods can also be applied to discrete location spaces by rounding each perturbed location to the nearest discrete point. In this case, they are more efficient than existing discrete methods. Trajectory utility of TraCS is theoretically analyzed and empirically evaluated. Evaluation results on discrete location spaces show that TraCS outperforms state-of-the-art methods in trajectory utility, particularly when the privacy parameter is large.
Ethical Considerations
This paper does not involve any human subjects or animals. While we use real-world trajectory data, they are publicly available, and all personal information has been anonymized.
Open Science
All code for the trajectory collection methods, evaluations, and datasets used in this paper will be made publicly available upon publication. We will include scripts for reproducing the experiments and generating the figures presented in the paper. An anonymized snapshot for review is currently available at https://anonymous.4open.science/r/TraCS-USENIX/README.md.
@@ Appendix A Related Work
This section reviews related trajectory collection methods under LDP and other quantifiable privacy notions. More privacy-preserving trajectory collection methods can be found in recent surveys [26, 3].
@@
We have disscussed three state-of-the-art trajectory collection methods under pure LDP: NGram [9], L-SRR [29], and ATP [37]. Among them, L-SRR was originally applied for origin-destination location pairs, while NGram, ATP, and this paper focus on trajectory-level perturbation. NGram attempts to ensure pure LDP for the “trajectory space”, also named as trajectory-level or instance-level LDP in surveys [26, 3], which is a stronger privacy than the location space. However, this is proved extremely challenging as the trajectory space increases exponentially with the number of possible locations in the location space. L-SRR, ATP and this paper focus on ensuring pure LDP for the location space. Unlike L-SRR, ATP and other existing methods based on discrete mechanisms, this paper focuses on continuous location spaces, which are stronger than any fine-grained discrete location space.
Another line of research is trajectory synthesis for privacy-preserving trajectory publication. A typical work is LDPTrace [10], which synthesizes trajectories using a generative model. LDPTrace discretizes the location space into cells and represents each trajectory as a sequence of cells. Transition patterns between neighboring cells are perturbed using LDP mechanisms and collected by an untrusted curator. The curator then trains a Markov chain model on the perturbed transition patterns to capture movement behaviors. Synthetic trajectories are generated from the trained model for publication. However, compared to methods that perturb each trajectory, such as NGram, L-SRR, and ATP, synthetic trajectories may be overly too random and not reflect the original trajectories.
@@
As a weaker privacy notion than LDP, (central) DP [12, 13] assumes a trusted curator who processes the raw data and releases the perturbed data. It requires that only neighboring locations in the location space be indistinguishable, rather than any two arbitrary locations. Compared to LDP, the concept of “neighboring” in DP has many interpretations. A comprehensive survey on DP-based trajectory collection and publication is available in [26].
The usage of direction information in trajectory collection dates back to an early work SDD [17], which attempts to add unbounded Laplace noise to the direction information. It provides a DP guarantee for unbounded continuous location spaces, but results in poor utility. The notion of “n-gram” in trajectory collection dates back to [5], where it abstracts a trajectory pattern as a gram. The idea of hierarchical decomposition can also be found in [36], where a spatial decomposition method is used for trajectory synthesis.
External knowledge is also widely used in DP-based trajectory collection methods. NGram’s prior work [8] provides a DP-based solution for trajectory synthesis by exploiting the public knowledge of the road network. Focusing on particular privacy-concerned locations, or sensitive zones [30], can also improve the trajectory utility while maintaining partial privacy guarantees.
@@
A widely used relaxed privacy notion than LDP for trajectory collection is Geo-indistinguishability [2, 25], where the indistinguishability between two locations is a function of their distance, i.e. privacy can decrease with distance. Input-discriminative LDP [14] goes one step further, assigning each location its own privacy level. Threshold-integrated LDP [35] defines a threshold to constrain the sensitive region near a location. This is similar to NGram’s reachability constraint but with a formal privacy definition. More relaxed LDP notions for trajectory collection can be found in paper survey [26]. Although these relaxed LDP notions generally provide better trajectory utility than pure LDP, they sacrifice the strong privacy guarantee of pure LDP, and often need hyperparameters to define the privacy level.
@@ Appendix B Proofs
Table 3: Notations used in this paper.
Notation Description
$\mathcal{M}_{\circ}$ Direction perturbation mechanism
$\mathcal{M}_{-}$ Distance perturbation mechanism
$\varphi^{\prime}$ Perturbed direction
$r^{\prime}(\varphi)$ Perturbed distance along $\varphi$
Table 3 summarizes the notations used throughout the remainder of the Appendix.
B.1 LDP Proof of Definition 5
@@ B.3 Proof of Theorem 3
The worst-case MSE is achieved when $\overline{r}=0$ or $\overline{r}=1$, i.e. at the endpoints (similar to the variance calculation in [31]). In this case,
@@ B.4 Proof of Theorem 4
Denote TraCS-C in Algorithm 2 as $\mathcal{M}$. We need to prove that $\mathcal{M}$ satisfies $\varepsilon$-LDP for each location $\tau\in\mathcal{S}$. In TraCS-C, each location $\tau$ is represented in Cartesian space $\mathcal{D}_{a}\times\mathcal{D}_{b}$ with unique coordinates $(a,b)$. Therefore, the key insight is to prove that the perturbation of $a$ and $b$ satisfies $\varepsilon/2$-LDP each.
In Theorem 2, we have already proved that $\mathcal{M}_{-}(\overline{a},\varepsilon/2)$ satisfies $\varepsilon/2$-LDP for the normalized distance $\overline{a}$. Then, the linear mapping from $\overline{a}\in[0,1)$ to $a\in\mathcal{D}_{a}$ is a post-processing step that does not affect the LDP property. This means $\mathcal{M}_{-}$ ensures $\varepsilon/2$-LDP for the distance space $\mathcal{D}_{a}$. In the same way, it also ensures $\varepsilon/2$-LDP for the distance space $\mathcal{D}_{b}$. Therefore, $\mathcal{M}=(\mathcal{M}_{-},\mathcal{M}_{-})$ satisfies $\varepsilon$-LDP for each location $\tau=(a,b)\in\mathcal{S}$ by Sequential Composition Theorem 1. (Alternatively, it can also be proved by directly computing the 2D pdfpdf ratio for obtaining the same location via $\mathcal{M}=(\mathcal{M}_{-},\mathcal{M}_{-})$.) Therefore, the entire trajectory with $n$ locations satisfies $n\varepsilon$-LDP. ∎
@@ Appendix C Complementary Materials
Figure 10: The Exponential mechanism $\mathcal{M}_{\text{exp}}(0,0)$ with $\varepsilon=4$ on two discrete location spaces within $0,1]\times[0,1]$. The number and arrangement of points affect the mechanism.
@@
Figure 11: TraCS-C mechanism $\mathcal{M}_{\text{TraCS-C}}(0,0)$ with $\varepsilon=4$ on $\mathcal{S}=[0,1]\times[0,1]$, which encompasses both discrete location spaces in Figure 10. The sampling probability is defined over areas rather than individual points.
C.1 Weaknesses of the Exponential Mechanism (Section 2.3)
C.1.1 High Computational Complexity
The biggest weakness of the Exponential mechanism is the complexity in calculating the score function and sampling from the probability distribution. Before perturbing location $x$, the Exponential mechanism requires computing the score function $d(x,y)$ for every possible location $y$ in the location space. If the location space has $m$ locations, the time complexity of computing the score function for a single location is $\mathcal{O}(m)$, which is computationally expensive for large-scale spaces. Moreover, ATP’s dynamic strategy to constrain the location space for perturbation makes it need to recompute the score function for each location, resulting in $\mathcal{O}(m)$ time complexity, where $m$ is the size of the location space.
In contrast, we will see that piecewise-based perturbation mechanisms in TraCS do not require score functions, and they sample from a $3$-piece CDF at each perturbation, resulting in a negligible constant time complexity, i.e. $\Theta(1)$, for each perturbation.
C.1.2 Affected by the Number and Arrangement of Locations
@@
C.2 Dominant Sector Comparison Between the Strawman Approach and TraCS-D (Example 1)

### v3

@@ 4.2.4. Time Cost Comparison
4.2.5. Experimental Results for Other Metrics
This subsection evaluates the performance of TraCS and existing methods under two additional metrics: range query preservation (Formula 17 in (DBLP:journals/pvldb/CunninghamCFS21)) and hotspot preservation (Section 6.2.4 in (DBLP:journals/pvldb/Zhang [^0] H23)). Given a threshold $\delta$, a perturbed location is considered correct if it lies within a $\delta$ distance of the corresponding sensitive location. Formally, for a trajectory $\mathcal{T}$ and its perturbed trajectory $\mathcal{T}^{\prime}$, the range query preservation (RQP) is defined as:
$RQP(\mathcal{T},\mathcal{T}^{\prime})=\frac{1}{|\mathcal{T}|}\sum_{i=1}^{|\mathcal{T}|}\mathbf{1}\{\|\tau_{i}-\tau^{\prime}_{i}\|_{2}\leq\delta\}\cdot 100\%,$
where $\mathbf{1}\{\cdot\}$ is the indicator function. A higher RQP indicates better trajectory utility. Hotspot preservation measures how many hotspots from a given set remain after perturbation. Given a set of hotspots $\mathcal{H}$, the locations in $\mathcal{H}$ are considered preserved if their perturbed locations after rounding are also in $\mathcal{H}$. Formally, the count difference (CD) of a trajectory in hotspot preservation is defined as:
$CD(\mathcal{T},\mathcal{T}^{\prime})=\sum_{i=1}^{|\mathcal{T}|}\mathbf{1}\{\tau_{i}\in\mathcal{H}\}-\sum_{i=1}^{|\mathcal{T}|}\mathbf{1}\{\tau_{i}\in\mathcal{H}\land\tau^{\prime}_{i}\in\mathcal{H}\},$
where $\mathbf{1}\{\cdot\}$ is the indicator function. A smaller CD indicates better trajectory utility. We can find the relationship between AE and these two metrics. For range query preservation with a given threshold $\delta$, when $AE\leq\delta$, the perturbed location is expected to satisfy the range query. For hotspot preservation a smaller AE leads to better hotspot retention after rounding in TraCS.
(a) Range query preservation
(b) Hotspot preservation
Figure 10. Comparison of other trajectory metrics on the CHI dataset. Higher RQP and lower ACD indicate better performance.
  • Range query preservation: We set the threshold $\delta=0.1$, and compute the RQP averaged over all trajectories.
  • Hotspot preservation: The hotspot set $\mathcal{H}$ is defined as the first $20\%$ frequent locations in the CHI dataset and we compute the CD averaged over all trajectories.
The results are shown in Figure 10. It can be seen that TraCS generally outperforms existing methods across different privacy levels, demonstrating better effectiveness in preserving both range queries and hotspots. Statistically, (i) the mean RQP across all $\varepsilon$ values for L-SRR, ATP, and NGram are 47.1%, 46.8%, and 56.5%, respectively, while the mean RQP of TraCS-D and TraCS-C are 58.6% and 68.4%, respectively. (ii) The mean CD across all $\varepsilon$ values for L-SRR, ATP, and NGram are 5.4, 4.9, and 5.2, respectively, while the mean CD of TraCS-D and TraCS-C are both 4.1. Particularly, TraCS shows improvement in hotspot preservation compared to existing methods across all privacy levels.
4.2.6. Assigning $\varepsilon$ for a Whole Trajectory
We also evaluate TraCS when the privacy parameter $\varepsilon$ is specified for an entire trajectory rather than per location. Specifically, for TraCS, L-SRR, NGram, and ATP, we set $\varepsilon$ at the trajectory level and allocate it uniformly across locations by assigning each location a parameter of $\varepsilon/|\mathcal{T}|$, where $|\mathcal{T}|$ is the trajectory length. This uniform allocation is simple and requires no additional information, and we consider it the most robust choice. In contrast, non-uniform allocations (e.g. assigning a larger parameter to utility-critical locations) may improve utility, but they typically rely on public notions of “importance” and can weaken privacy protection for those locations. Appendix C.10 provides results and discussion on this evaluation, showing that TraCS outperforms existing methods in trajectory utility under this setting when $\varepsilon$ is large.
4.3. Summary of Evaluation
@@
These findings validate the effectiveness and efficiency of TraCS in collecting trajectory data under LDP in both continuous and discrete spaces.
5. Conclusions
This paper presents TraCS-D and TraCS-C, providing local differential privacy guarantees for trajectory collection in continuous location spaces. TraCS-D uses a novel direction-distance perturbation procedure, while TraCS-C perturbs the Cartesian coordinates of each location. These two methods can also be applied to discrete location spaces by rounding each perturbed location to the nearest discrete point, achieving better efficiency than existing discrete methods. Trajectory utility of TraCS is analyzed theoretically and evaluated empirically. Evaluation results on discrete location spaces show that TraCS outperforms state-of-the-art discrete methods in trajectory utility, particularly when the privacy parameter is large.
In summary, TraCS demonstrates that collecting trajectory data in continuous spaces generally better discrete approaches, avoiding privacy and efficiency issues caused by discretization, while can also be rounded to discrete spaces when necessary.
@@ Appendix A Related Work
This section reviews related trajectory collection methods under LDP and other quantifiable privacy notions. More privacy-preserving trajectory collection methods can be found in recent surveys (DBLP:journals/popets/MirandaPascualGPFS23; DBLP:journals/popets/BuchholzAWNK24).
@@
We have disscussed three state-of-the-art trajectory collection methods under pure LDP: NGram (DBLP:journals/pvldb/CunninghamCFS21), L-SRR (DBLP:conf/ccs/WangH0QH22), and ATP (DBLP:journals/pvldb/Zhang [^0] H23). Among them, L-SRR was originally applied for origin-destination location pairs, while NGram, ATP, and this paper focus on trajectory perturbation. NGram attempts to ensure pure LDP for the “trajectory space”, also named as trajectory-level or instance-level LDP in surveys (DBLP:journals/popets/MirandaPascualGPFS23; DBLP:journals/popets/BuchholzAWNK24), which is a stronger privacy than the location space. However, this is proved extremely challenging as the trajectory space increases exponentially with the number of possible locations in the location space. L-SRR, ATP and this paper focus on ensuring pure LDP for the location space. Unlike L-SRR, ATP and other existing methods based on discrete mechanisms, this paper focuses on continuous location spaces, which are more expressive than any fine-grained discrete location space.
Beyond the above three methods that design LDP mechanisms for trajectory collection, several works leverage external knowledge to improve the utility of LDP-based trajectory collection. WF-LDPSR (DBLP:journals/compsec/LiXZZ25) adaptively determines each user’s privacy protection level based on the sensitivity of the user’s information, and applies a water-filling strategy to allocate privacy budgets during perturbation. Regional popularity (i.e. the popularity of different regions in the location space) can also be exploited to constrain trajectory boundaries and improve utility (DBLP:journals/tifs/ZhangNFHZ25). In addition, recent work studies poisoning attacks against NGram and ATP (DBLP:journals/corr/abs-2503-07483), where adversaries induce malicious users to submit poisoned trajectories to the data collector with the goal of promoting target patterns.
Another line of research is trajectory synthesis for privacy-preserving trajectory publication. A typical work is LDPTrace (DBLP:journals/pvldb/DuHZFCZG23), which synthesizes trajectories using a generative model. LDPTrace discretizes the location space into cells and represents each trajectory as a sequence of cells. Transition patterns between neighboring cells are perturbed using LDP mechanisms and collected by an untrusted curator. The curator then trains a Markov chain model on the perturbed transition patterns to capture movement behaviors. Following LDPTrace, ADGTrace (DBLP:journals/tmc/CaiLSZYZX26) also employs Markov chain models for trajectory synthesis; it does not provide LDP guarantees and leverages personal features to improve utility. Compared to methods that perturb each trajectory, such as NGram, L-SRR, and ATP, synthetic trajectories may be overly too random and not reflect the original trajectories if not personalized enough.
LDP Mechanisms for Bounded Numerical Domains
The perturbation mechanisms used in TraCS for direction and distance perturbation are utility-optimized piecewise-based mechanisms in OGPM (DBLP:journals/popets/ZhengMH25). Other LDP mechanisms for bounded numerical domains (DBLP:conf/ccs/WeggenmannK21; DBLP:journals/tdp/000619; DBLP:journals/jpc/HolohanABA20) can also be incorporated into TraCS; refer to Section 3.5.6 for discussions and comparisons. Compared with OGPM, which focuses on optimizing the 1D mechanisms’ data utility, the design of 2D mechanisms in TraCS needs to consider the unique characteristics of trajectory data and 2D spaces. In particular, the direction-distance perturbation in TraCS-D to guarantee LDP for a rectangular region requires careful design. Meanwhile, this paper emphasizes the overlooked fundamental issues inherent in discrete-space LDP mechanisms for trajectory collection (DBLP:journals/pvldb/CunninghamCFS21; DBLP:journals/pvldb/Zhang [^0] H23; DBLP:conf/ccs/WangH0QH22), which make privacy, utility, and efficiency discretization-dependent and prevent guarantees for continuous spaces. TraCS demonstrates that designing mechanisms directly in continuous space (via building blocks like OGPM, etc.) avoids these issues and subsumes discrete cases via rounding.
@@
As a weaker privacy notion than LDP, (central) DP (DBLP:conf/icalp/Dwork06; DBLP:journals/fttcs/DworkR14) assumes a trusted curator who processes the raw data and releases the perturbed data. It requires that only neighboring locations in the location space be indistinguishable, rather than any two arbitrary locations. Compared to LDP, the concept of “neighboring” in DP has many interpretations. A comprehensive survey on DP-based trajectory collection and publication is available in (DBLP:journals/popets/MirandaPascualGPFS23).
The use of direction information in trajectory collection dates back to an early work SDD (DBLP:conf/ssdbm/JiangSBKT13), which attempts to add unbounded Laplace noise to the direction information. It provides a DP guarantee for unbounded continuous location spaces, but results in poor utility. The notion of “n-gram” in trajectory collection dates back to (DBLP:conf/ccs/ChenAC12), where it abstracts a trajectory pattern as a gram. The idea of hierarchical decomposition can also be found in (DBLP:conf/sigmod/ZhangXX16), where a spatial decomposition method is used for trajectory synthesis.
External knowledge is also widely used in DP-based trajectory collection methods. NGram’s prior work (DBLP:conf/ssd/CunninghamCF21) provides a DP-based solution for trajectory synthesis by exploiting the public knowledge of the road network. Focusing on particular privacy-concerned locations, or sensitive zones (DBLP:journals/compsec/WangK20), can also improve the trajectory utility while maintaining partial privacy guarantees.
@@
A widely used relaxed privacy notion than LDP for trajectory collection is Geo-indistinguishability (DBLP:conf/ccs/AndresBCP13; DBLP:journals/popets/MendesCV20), where the indistinguishability between two locations is a function of their distance, i.e. privacy can decrease with distance. Input-discriminative LDP (DBLP:conf/icde/Gu0XC20) goes one step further, assigning each location its own privacy level. Threshold-integrated LDP (10.1007/978-981-97-5562-2_5) defines a threshold to constrain the sensitive region near a location. This is similar to NGram’s reachability constraint but with a formal privacy definition. More relaxed LDP notions for trajectory collection can be found in paper survey (DBLP:journals/popets/MirandaPascualGPFS23). Although these relaxed LDP notions generally provide better trajectory utility than pure LDP, they sacrifice the strong privacy guarantee of pure LDP, and often need hyperparameters to define the privacy level.
@@ Appendix B Proofs
B.1. LDP Proof of Definition 1
@@ B.3. Proof of Theorem 5
The worst-case MSE is achieved when $\overline{r}=0$ or $\overline{r}=1$, i.e. at the endpoints (similar to the variance calculation in (DBLP:conf/icde/WangXYZHSS019)). In this case,
@@ B.4. Proof of Theorem 6
Denote TraCS-C in Algorithm 2 as $\mathcal{M}$. We show that $\mathcal{M}$ satisfies $\varepsilon$-LDP for each location $\tau\in\mathcal{S}$. In TraCS-C, each location $\tau$ is represented in the Cartesian space $\mathcal{D}_{a}\times\mathcal{D}_{b}$ with unique coordinates $(a,b)$. Thus, it suffices to prove that the perturbations of $a$ and $b$ each satisfy $\varepsilon/2$-LDP.
From Theorem 4, $\mathcal{M}_{-}(\overline{a};\varepsilon/2)$ satisfies $\varepsilon/2$-LDP for the normalized coordinate $\overline{a}\in[0,1)$. The subsequent linear mapping from $\overline{a}\in[0,1)$ to $a\in\mathcal{D}_{a}$ is post-processing and therefore preserves the LDP guarantee. Hence, $\mathcal{M}_{-}$ ensures $\varepsilon/2$-LDP over $\mathcal{D}_{a}$. By the same argument, $\mathcal{M}_{-}$ also ensures $\varepsilon/2$-LDP over $\mathcal{D}_{b}$.
Therefore, $\mathcal{M}\coloneqq(\mathcal{M}_{-},\mathcal{M}_{-})$ satisfies $\varepsilon$-LDP for each location $\tau=(a,b)\in\mathcal{S}$ by the Sequential Composition Theorem 2. Equivalently, one can verify the guarantee by directly bounding the 2D pdfpdf ratio induced by $\mathcal{M}$. Consequently, the entire trajectory with $n$ locations satisfies $n\varepsilon$-LDP. ∎
@@ Appendix C Complementary Materials
C.1. Space-Dependence of Indistinguishability (Section 1)
Indistinguishability of a data point is always relative to a specified data space, which determines the set of alternatives that an adversary may try to distinguish it from. Accordingly, indistinguishability cannot be meaningfully discussed without first specifying the underlying space, for two main reasons.
(i) LDP perspective. The LDP guarantee is defined with respect to a particular input domain. If two algorithms operate on the same domain and use the same privacy parameter $\varepsilon$, then their privacy guarantees are comparable (in the sense of the LDP definition); if their domains differ, then their guarantees may not be directly comparable, even when they share the same $\varepsilon$.
(ii) Information-theoretic perspective. For an $\varepsilon$-LDP mechanism $\mathcal{M}$, the mutual information $I(x;\mathcal{M}(x))$ is upper bounded by a function of both $\varepsilon$ and the size of the input space, i.e. $I(x;\mathcal{M}(x))\leq\mathcal{O}(\varepsilon,|\mathcal{X}|)$, where $|\mathcal{X}|$ denotes the cardinality of the input space $\mathcal{X}$. A similar bound can be found in (Duchi2013LocalPD). [^1] Intuitively, LDP enforces pairwise indistinguishability but does not directly account for how indistinguishability accumulates over many alternatives, which can be captured by mutual information.
This observation is particularly relevant to trajectory collection or synthesis under LDP when discretization-based methods are used. Even for a fixed continuous area, different discretization strategies (e.g. uniform grids, adaptive grids, or different sets of points of interest (POIs)) induce different discrete location spaces with different sizes and spatial layouts. Consequently, the effective indistinguishability provided by the same LDP mechanism can vary substantially across discretizations, even when the same $\varepsilon$ is used.
C.2. Limitations of the Exponential Mechanism (Section 1 and Section 2.3)
C.2.1. High Computational Complexity
The biggest limitation of the Exponential mechanism is the complexity in calculating the score function and sampling from the probability distribution. Before perturbing location $x$, the Exponential mechanism requires computing the score function $d(x,y)$ for every possible location $y$ in the location space. If the location space has $m$ locations, the time complexity of computing the score function for a single location is $\mathcal{O}(m)$, which is computationally expensive for large-scale spaces. Moreover, ATP’s dynamic strategy to constrain the location space for perturbation makes it need to recompute the score function for each location, resulting in $\mathcal{O}(m)$ time complexity, where $m$ is the size of the location space.
In contrast, we have seen that piecewise-based perturbation mechanisms in TraCS do not require score functions, and they sample from a $3$-piece CDF at each perturbation, resulting in a negligible constant time complexity, i.e. $\Theta(1)$, for each perturbation.
@@ C.2.1. High Computational Complexity
Figure 11. The Exponential mechanism $\mathcal{M}_{\text{exp}}(0,0)$ with $\varepsilon=4$ on two discrete location spaces within $0,1]\times[0,1]$. The number and arrangement of points affect the mechanism.
@@
Figure 12. TraCS-C mechanism $\mathcal{M}_{\text{TraCS-C}}(0,0)$ with $\varepsilon=4$ on $\mathcal{S}=[0,1]\times[0,1]$, which encompasses both discrete location spaces in Figure 11. The sampling probability is defined over areas rather than individual points.
C.2.2. Affected by the Number and Arrangement of Locations
@@
C.3. Limitations of Applying Discrete LDP Mechanisms to Continuous Spaces (Section 1)
Although a continuous space can be discretized before applying discrete LDP mechanisms, this approach inherits the privacy–utility–efficiency issues of discrete methods and introduces additional challenges in choosing an appropriate discretization strategy.
Gridding the continuous space. A common discretization approach is to partition the continuous space into uniform grid cells and treat each cell as a discrete input and output of the discrete mechanism. The reported cell can then be post-processed by randomly sampling a location within that cell.
(a) Error of the gridding approach with different cell sizes.
(b) Average time cost of perturbing one location.
Figure 13. Performance of the gridding approach with different cell sizes in a continuous space. Coarse-grained grids (e.g. $10\times 10$) perform better when $\varepsilon$ is small, whereas fine-grained grids (e.g. $100\times 100$) perform better when $\varepsilon$ is large, at the cost of higher computational time. In contrast, TraCS achieves better utility than both gridding baselines across all $\varepsilon$ values, with negligible time overhead.
Dilemma. However, the gridding approach faces a fundamental dilemma in choosing the grid cell size, which directly affects the privacy–utility–efficiency tradeoff. (i) If the grid cells are too large (i.e. coarse-grained gridding), the mechanism is more likely to output the true cell, but the intra-cell error (i.e. the distance between the true location and the post-processed location sampled within the same cell) can be large. (ii) If the grid cells are too small (i.e. fine-grained gridding), the intra-cell error can be reduced, but the probability of outputting the true cell decreases. Meanwhile, the computational cost increases with the number of cells.
We empirically evaluate the gridding approach under different cell sizes in a continuous space and compare it with TraCS-C. Specifically, we consider the $0,1]\times[0,1]$ location space and two grid resolutions: $10\times 10$ and $100\times 100$. For a fixed location, we generate $500$ perturbed locations using (i) the gridding approach with the Exponential mechanism (EM), where the score is defined by the distance between cell centers, and (ii) TraCS-C. We then report the average error and the average time required to perturb a single location. Figure 13 summarizes the results. In the small-$\varepsilon$ regime, coarse-grained grids (e.g. $10\times 10$) outperform fine-grained grids (e.g. $100\times 100$), whereas in the large-$\varepsilon$ regime, fine-grained grids perform better. This trend is consistent with the above dilemma: it is difficult to select a single grid resolution that performs well across different $\varepsilon$ values. In contrast, TraCS-C consistently achieves better utility than both gridding baselines for all tested $\varepsilon$ values, while incurring negligible time overhead.
C.4. Dominant Sector Comparison Between the Strawman Approach and TraCS-D (Example 2)